OpenStudy (anonymous):

-7T-20+7=-7-3T solve the equation

OpenStudy (anonymous):

-7t - 13 = -7 - 3t -7t + 3t = -7 + 13 -4t = 6 t = -6/4 = -3/2

OpenStudy (anonymous):

\[T=-6/4= -3/2\]\[4T=-6\]\[-7T-13=-7-3T\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are most welcome....:)